OpenStudy (kaylaprincess):

is point slope and slope intercept the same? both end with an answer in the style of : y = mx + b

OpenStudy (anonymous):

Slope intercept is the equation, slope is rise over run

OpenStudy (jadeishere):

Technically, no they're not the same. Would you care to know why?

OpenStudy (kaylaprincess):

yes of course cx

OpenStudy (jadeishere):

Lol okay. They have different formulas, basically. slope-intercept form is where y = mx + b point-slope form is where \[y- y _{1}= m(x - x _{1})\] you are right that they end in slope-intercept, but most often, most linear equations do :) so i hope this helps >.<

OpenStudy (kaylaprincess):

thanks man http://img.pandawhale.com/80419-dwight-schrute-fist-pump-screa-CtKy.gif

OpenStudy (jadeishere):

LOL!! You're welcome :D